OpenStudy (anonymous):

Find the diagonal of a square whose sides are of the given measure. Given = 5" *the answer has a 2 on it *

OpenStudy (anonymous):

diagonal = sqr(2) * 5

OpenStudy (anonymous):

as D= sqr(2) * a........for square (a=b)